A lake shore lot was described in a real estate magazine.
Its frontage is 100 FEET.
Its depth is 90 FEET and its lake front is 150 FEET.
How many square FEET does this lot contains?

Answer by stanbon(75887)

A lake shore lot was described in a real estate magazine.
Its frontage is 100 FEET.
Its depth is 90 FEET and its lake front is 150 FEET.
How many square FEET does this lot contains?
-------------------
Area of a trapezoid = (1/2)h(b1 + b2)
---------------------
A = (1/2)90(100+150)
A = 45*250
A = 11,250 sq. ft.
=====================
Cheers,
Stan H.